Anle Corporation has a current stock price of

$ 16.24$16.24

and is expected to pay a dividend of

$ 0.85$0.85

in one year. Its expected stock price right after paying that dividend is

$ 18.06$18.06.

a. What is​ Anle's equity cost of​ capital?

b. How much of​ Anle's equity cost of capital is expected to be satisfied by dividend yield and how much by capital​ gain?

a. What is​ Anle's equity cost of​ capital?

​Anle's equity cost of capital is

nothing​%.

​(Round to two decimal​ places.)

b. How much of​ Anle's equity cost of capital is expected to be satisfied by dividend yield and how much by capital​ gain?

The portion of​ Anle's equity cost of capital that is expected to be satisfied by the dividend yield is

nothing​%.

​(Round to two decimal​ places.)The portion of​ Anle's equity cost of capital that is expected to be satisfied by capital gains is

nothing​%.

​(Round to two decimal​ places.)
Answer #1
a] Equity cost of capital = (0.85+18.06)/16.24-1 = 16.44%
b] Dividend yield = 0.85/16.24 = 5.23%
Capital gain yield = 18.06/16.24-1 = 11.21%